# Rise and fall of nonstabilizerness via random measurements

### v1

@@ Appendix D Deviation from Haar-random behavior
Figure 15: Protocol with $\theta_{M}=0$. a) $\mathcal{M}_{2}$ against $\nu$ for initial Haar-random states for different $N$, plotted together with the analytic prediction in grey dotted line. b) $\mathcal{M}_{2}$ against $\nu$ for initial $|T\rangle^{\otimes N}$ for different $N$, plotted together with the analytic prediction in grey dotted line. c) $\mathcal{M}_{2}$ against $\nu$ for initial $|\psi\rangle_{\text{GUE}}$ state for different $N$. Inset: zoom on main plot. In this case the legend was remove for clarity.
@@ Appendix E Qubit number dependence of steady-state SREs
We now study how the steady-state of our circuit depending on qubit number $N$. For $\theta_{M}=0$, the steady-state is trivially zero as the final state is a random stabilizer state. In particular, in Fig. 16 we show the steady-state SRE $\mathcal{M}_{2}^{\text{ss}}$ against $N$ for $\theta_{M}=1$. We find as expected that asymptotically, the SRE increases linearly with $N$. As reference, we also plot the average SRE of Haar-random states for $N-2$ qubits. Curiously, we find that the steady-state SRE for $N$ qubits nearly matches the SRE of Haar random states of $N-2$. While the steady-state does not have the characteristics of a Haar random state, we leave as an open question whether this relationship between steady-state and Haar random magic has a deeper physical meaning.
@@
Figure 16: Steady-state SRE ¯$\overline{\mathcal{M}}_{2}$ against $N$ with $\theta_{M}=1$ are shown as dots, while the average SRE of Haar-random states of $N-2$ qubits as solid line.

### v2

@@ Appendix D Deviation from Haar-random behavior
Figure 15: Protocol with $\theta_{M}=0$. a) $\mathcal{M}_{2}$ against $\nu$ for initial Haar-random states for different NN, plotted together with the analytic prediction in grey dotted line. b) $\mathcal{M}_{2}$ against $\nu$ for initial $|T\rangle^{\otimes N}$ for different NN, plotted together with the analytic prediction in grey dotted line. c) $\mathcal{M}_{2}$ against $\nu$ for initial $|\psi\rangle_{\text{GUE}}$ state for different NN. Inset: zoom on main plot. In this case the legend was remove for clarity.
@@ Appendix E Qubit number dependence of steady-state SREs
We now study how the steady-state of our circuit depending on qubit number $N$. For $\theta_{M}=0$, the steady-state is trivially zero as the final state is a random stabilizer state. In particular, in Fig. 16 we show the steady-state SRE $\mathcal{M}_{2}^{\text{ss}}$ against $N$ for $\theta_{M}=1$. We find as expected that asymptotically, the SRE increases linearly with $N$. As reference, we also plot the average SRE of Haar-random states for N2N-2 qubits. Curiously, we find that the steady-state SRE for $N$ qubits nearly matches the SRE of Haar random states of N2N-2. While the steady-state does not have the characteristics of a Haar random state, we leave as an open question whether this relationship between steady-state and Haar random magic has a deeper physical meaning.
@@
Figure 16: Steady-state SRE $\overline{\mathcal{M}}_{2}$ against NN with $\theta_{M}=1$ are shown as dots, while the average SRE of Haar-random states of N2N-2 qubits as solid line.